OpenStudy (anonymous):

Monthly = P (1 + r/12)12 = (monthly compounding) Karan starts saving for a down-payment on a house by making regular payments at the end of each month into an account that earns 8.4% / a interest, compounded monthly. Determine the amount of each monthly payment needed to accumulate $10 000 in three years. @terenzreignz

terenzreignz (terenzreignz):

Let's see...

terenzreignz (terenzreignz):

Okay, so r = 8.4% = 0.084 right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

But then again, regular monthly payments are made... looks like this is an annuity. Am I right, @UnkleRhaukus ? (sorry, money isn't really my thing :D )

OpenStudy (anonymous):

yeah its annuity, its not my thing either aha

terenzreignz (terenzreignz):

Okay, so we have a formula for an annuity \[\huge a_{{n}}=\frac{1-(1+i)^{-n}}{i}\] right?

terenzreignz (terenzreignz):

Where i is the effective rate of interest per period.

terenzreignz (terenzreignz):

Catch me so far?

terenzreignz (terenzreignz):

Sorry, \(\large a_n\) is the present value of an annuity that pays 1 per period.

OpenStudy (anonymous):

yeah I'm following

terenzreignz (terenzreignz):

But then again, we are given the accumulated value... \[\huge s_n = \frac{(1+i)^n-1}{i}\]

terenzreignz (terenzreignz):

So... we need to find how much she puts in every month, right?

OpenStudy (anonymous):

yes that is correct

terenzreignz (terenzreignz):

Okay, the accumulated value at the end of n-periods of a deposit of x per period is given by \[\huge x \cdot s_n\]

terenzreignz (terenzreignz):

Now, it is given that \[\huge x\cdot s_n = 10000\]

terenzreignz (terenzreignz):

Let's take this per month, shall we?

terenzreignz (terenzreignz):

After all, she is depositing on a monthly basis.

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

so... if it's 3 years, how many months is that?

OpenStudy (anonymous):

sorry, 36

OpenStudy (anonymous):

n=36 right?

terenzreignz (terenzreignz):

yeah, 36. Now what's the effective rate of interest per month?

OpenStudy (anonymous):

0.084

terenzreignz (terenzreignz):

No... that's for 12 months, remember? :P

OpenStudy (anonymous):

hmm

terenzreignz (terenzreignz):

the interest per month is 1/12 of the annual compounded interest :)

OpenStudy (anonymous):

do i use r=e-1 to find it

OpenStudy (anonymous):

i'm lost

terenzreignz (terenzreignz):

No, the interest per month is simply \[\huge \frac{0.084}{12}=0.007\]

OpenStudy (anonymous):

thats what i did originally but i confused myself ...

terenzreignz (terenzreignz):

well, don't... \[\huge x \cdot s_n = 10000\] \[\huge s_n = \frac{(1+i)^n-1}{i}\] \[\huge i=0.007\]\[\huge n =36 \ (months)\] Get to it, champ :)

OpenStudy (anonymous):

aha thank you :)

OpenStudy (anonymous):

10 000=(1+0.007)36-1 10 000= (1.007)36-1 10 000= 35.25 -1 ? am i doing this right so far?

terenzreignz (terenzreignz):

you forgot to divide by i And where's the x?

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

k i confused myself

OpenStudy (anonymous):

10 000 = 35.25/0.007 ?

terenzreignz (terenzreignz):

Be careful with your applications of the operations... What is \[\huge 1.007^{36}\]?

OpenStudy (anonymous):

1.007 to the power of 36

terenzreignz (terenzreignz):

yeah, evaluate it... using your calculator.

OpenStudy (anonymous):

1.28547

terenzreignz (terenzreignz):

And use that.

terenzreignz (terenzreignz):

\[\huge s_n = \frac{(1+i)^n-1}{i}\]

terenzreignz (terenzreignz):

I have to go now, ok? Good luck, sorry I couldn't stay longer... ----------------------------------------------- Terence out

OpenStudy (anonymous):

thank you so much for your help !